What is the mass of 3.40 x 10^22 molecules of glucose? C6H12O6


1
Expert's answer
2020-12-18T08:43:39-0500

Molar mass of glucose , C6H12O6

= [(6×12)+(12×1)+(6×16)] g/mol

= 180 g/mol


6.023×10^23 molicules = 1 mole

So,Mass of

6.023×10^23 molicules of glucose = 180g

3.40 x 10^22 molicules of glucose = [3.40 x 10^22/(6.023×10^23)]×180 g

= 10.16 g


Hence, mass of 3.40 x 10^22 molecules of glucose = 10.16 g